OpenStudy (anonymous):

Can someone help me simplify the following radicands? The problem is in the comments! Thank you!

OpenStudy (anonymous):

\[3\sqrt{2x} + x \sqrt{8x} - 5\sqrt{18x}\]

OpenStudy (anonymous):